# Improving Relational Database Interactions with Large Language Models: Column Descriptions and Their Impact on Text-to-SQL Performance

### v1

@@ 3.2 Improving Existing Column Descriptions with LLMs
Easy 3.83 3.55 3.26 2.98 2.97 3.24
Medium 3.66 3.28 3.18 2.89 2.79 2.84
Hard 3.46 2.46 2.90 2.76 2.69 2.04
Very Hard 1.93 1.24 1.59 1.47 1.44 1.12
@@
Table 3: Mean quality scores for each model’s generated descriptions across the column difficulty levels in the ColSQL dataset, as rated by the human evaluators.
3.3 Creating the Gold Standard Dataset
Since not all LLM-generated improvements were perfect, manual corrections of the generations were necessary to achieve the high standard required for a gold standard dataset. First, disagreements in scores were addressed by reviewing each other’s annotations and then through discussion to reach a consensus. Descriptions that did not achieve a “Perfect” quality score from both annotators after this phase were manually corrected until both annotators agreed on a “Perfect” rating. The final set of descriptions was then reviewed to ensure they provided clear, comprehensive, and contextually rich information for each column. Through this process, we believe we have created a dataset of high quality, serving as a reliable reference for evaluating the performance of models in generating column descriptions. The final composition of the dataset can be found in Table 1. Examples of gold descriptions can be found in Appendix B.3.
3.4 Annotating Column Difficulty
Following the finalization of the gold standard descriptions, each column was assessed on the difficulty in generating descriptions for it. Columns were categorized into four levels: "Easy," "Medium," "Hard," and "Very Hard." These levels reflect the annotators’ assessment of the amount of available information necessary to generate an accurate description for a column. Examples of columns of each difficulty can be found in B.3.
At the "Easy" level, annotators could determine the column description by only looking at the schema of the relevant table. For example, in the Financial database, the Client table’s birth_date column is self-explanatory.
At the "Medium" level, annotators could accurately determine the column description given example data in addition to the table schema from the database. This suggests that while some additional context is required, the column’s role can still be deduced with a reasonable amount of information. For instance, in the student_club database, the zip_code table contains a column named short_state. Given the data inside the database, it is possible to accurately determine that this column contains two-letter abbreviations of states.
Columns categorized as "Hard" present greater challenges. Even with the schema and example data from the database, annotators were unsure about what the column description should be. This uncertainty arises from ambiguities or complexities that cannot be resolved with the given information, highlighting the need for further context or domain knowledge. For example, in the formula_1 database’s results table, the rank column has data indicating rankings, but without context, it is hard to determine if it refers to race position, qualifying rank, or another metric. The description, "The overall rank of the driver’s fastest lap time in the race," clarifies its purpose.
The "Very Hard" level involved columns where accurate descriptions were impossible without additional documentation. An example is the A11 column in the Districts table of the Financial database, which contains only numbers and a completely uninformative name.
The assessment was performed by two independent annotators, distinct from those who created the gold standard descriptions, but with the same level of expertise. The inter-annotator agreement for difficulty level was fair, with a Cohen’s kappa of 0.31.
After completing the annotation process, the annotators used the same procedure as for the gold descriptions to resolve disagreements. It became clear during their discussion that one annotator had misunderstood the difference between the "Easy" and "Medium" criteria, causing the low agreement score. This annotator incorrectly thought "Easy" included only columns with unambiguous names, whereas our definition allows considering other column names in the table. All remaining disagreements were resolved through discussion, resulting in one final difficulty score for each column.
The distribution of columns across the four difficulty levels is presented in Table 1. We observed that a significant portion of columns fell into the “Easy” and “Medium” categories, with “Very Hard” columns primarily residing in the Financial domain, where tables often have completely uninformative column names such as “A1”.
4 Generating Column Descriptions
In this section, we assess LLMs’ ability to generate detailed column descriptions using our dataset, using both human evaluations and the LLM-as-a-judge.
Annotator/LLM Command R+ CodeStral 22B Mistral 7b Qwen 2 72B Mixtral 22B GPT-4o
Annotator 1 vs GPT-4o 0.22 0.16 0.24 0.21 0.16 0.12
Annotator 2 vs GPT-4o 0.25 0.18 0.23 0.22 0.17 0.12
Annotator 1 vs Annotator 2 0.60 0.71 0.71 0.62 0.63 0.64
Table 4: Cohen’s Kappa scores between annotators, and between annotators and a GPT-4o LLM judge.
4.1 Models
To evaluate the effectiveness of LLMs in generating column descriptions, we used our dataset presented in Section $3.$ We evaluated the following models: GPT-4o, Mixtral-8x22B-Instruct-v0.1 Jiang et al. (2024), Command R+, Qwen2-72B-Instruct Yang et al. (2024), Codestral-22B-v0.1, and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). These models were selected to provide a diverse range of perspectives, including a state-of-the-art closed-source model (GPT-4o), several newly released high-performing open-source models (Qwen2-72B, Mixtral-8x22B, Codestral-22B, Command R+), and a smaller model (Mistral-7B).
Inference for the open-source models was performed on several A100 GPUs, using BitsandBytes Dettmers et al. (2022) to load the models in 8-bit format. We set the model temperature to 0.7.
4.2 Human Evaluation
@@
For each model’s generated output, two human annotators independently assessed the quality of the column descriptions. As previously, the set of annotators was different from the ones that created the gold dataset, but with the same type of expertise. The annotations were also based on a slightly different set of criteria from what was used when rating the original and improved column descriptions, presented in Section 3.2, as we now have the gold column descriptions as a reference point.
@@
The generated outputs were scored for their usefulness, using labels “Perfect”, “Almost Perfect”, “Somewhat correct”, “Incorrect”, and “No Description”. For example, the main difference between “Perfect” and “Almost Perfect”, is that the latter contains additional redundant information that is not incorrect. A “Somewhat correct” column description is correct, but there is information missing. The full instructions and definitions of the labels are presented in Appendix B.1.4.
@@
To ensure the reliability of our evaluations, we measured the inter-annotator agreement for each model’s output. Cohen’s kappa for each model is presented in Table 4. The agreement among annotators was above 0.60 for all models, indicating a substantial agreement between the annotators.
The results of the model evaluations are presented in Figure 2 and Table 3. Our observations show that GPT-4o and Command R+ generated the highest proportion of "Perfect" descriptions, indicating their ability to provide concise and accurate column descriptions. In contrast, Qwen 2 72B and Codestral 22B had a high number of "Almost perfect" descriptions, suggesting a tendency to be verbose and add unnecessary information. Incorrect generations were relatively uncommon for most models, which might be more important than the distinctions between different levels of correctness. However, the ratio of incorrect answers were higher for smaller models like Mistral 7B.
4.3 Evaluating Models Using LLM-as-a-Judge
We also performed an evaluation using LLM-as-a-judge, with GPT-4o as the judge. This evaluation aimed to investigate whether there exists an automated metric that could effectively replace human evaluations.
We compared the model rankings as judged by GPT-4o to those determined by human annotators to see if there was a consistent correlation. The results, as shown in Table 4, indicate low alignment between GPT-4o and human annotators, with Cohen’s Kappa scores ranging from 0.12 to 0.25. Notably, the agreement between human annotators was significantly higher, with Kappa scores ranging from 0.60 to 0.71.
As a small side study, we also combined the “Perfect” and “Almost” Perfect’ scores, resulting in a doubling of Kappa values. Although the scores remain modest, this suggests potential for refining the evaluation process to better align with human ratings.
5 Impact of Column Descriptions for Text-to-SQL
To study the usefulness of column descriptions for text-to-SQL systems, we evaluated the best-performing model (GPT-4o) on the development set of BIRD-Bench, which contains 1534 natural language questions and SQL query pairs. Our goal was to isolate and understand the effects of column descriptions on model performance.
The experimental setup involved four scenarios:
  1. 1.
    Without any column descriptions: The model operates with only the database schema.
  2. 2.
    With BIRD-Bench original column descriptions: The model uses the original descriptions provided by BIRD-Bench in addition to the schema.
  3. 3.
    With generated column descriptions: The model uses descriptions generated by the best model from Section 4 in addition to the schema, simulating a realistic scenario.
  4. 4.
    With gold standard descriptions: The model uses the manually refined gold standard descriptions in addition to the schema, representing the best-case scenario.
Included Metadata Execution Accuracy
Without col. descr. 0.30
With BIRD-Bench col. descr. 0.31
With GPT-4o predicted col. descr. 0.32
With gold col. descr. 0.33
Table 5: Execution Accuracy over the BIRD-Bench development set for a zero-shot text-to-sql model using GPT-4o with and without different column descriptions.
In all scenarios, we included only column descriptions as the available metadata alongside the database schema. The full prompt for the text-to-SQL generation is presented in Appendix B.4.
The full results are presented in Table 5. Our findings indicate that column descriptions improve model performance on text-to-SQL tasks. We also observe that the GPT-4o generated column descriptions, which have been generated using only the database schema and example data, is slightly better than the existing BIRD-Bench column descriptions, which were created by human annotators.
We notice an improvement for the model with gold descriptions over the model with access to no column descriptions for questions involving difficult columns, where 34% of uniquely correct text-to-SQL samples referenced columns of a ”Hard“ or ”Very Hard“ difficulty rating. This shows that additional contextual information is crucial. In the next section we perform a more detailed study on how column descriptions can help models correctly predict text-to-SQL questions that reference “Very Hard” columns.
5.1 Analyzing the Effects of Difficult Columns
To better understand the usefulness of column descriptions in low-information scenarios, we replaced all the column names in each database with uninformative column names. The first table in database A where denoted by A1, the second by A2, and so on.
@@ 6 Conclusions and Future Work
In this study, we have created ColSQL, a dataset with curated human annotations of column descriptions. This dataset serves as a reliable gold standard for evaluating the performance of models in generating accurate and contextually rich column descriptions.
Our main findings demonstrate that LLMs can effectively generate high-quality column descriptions. Among the models evaluated, GPT-4o and Command R+ performed best, particularly in generating descriptions that meet our "Perfect" criteria where GPT-4o meets the criteria for 75% of the columns.
Through our evaluation process, supported by rigorous human annotations, we have highlighted both the strengths and limitations of various models in this task and shown that LLM-as-a-judge is not yet viable alternative to human annotations when ranking models for this task. We leave it to future work to find reliable automatic evaluations for this task.
We have also shown that detailed column descriptions significantly enhance text-to-SQL systems, especially for columns that are completely uninformative. For uninformative columns we see a 20% increase. This underscores the importance of metadata as a bridge between the database system and generative models interacting with that system.
Future work will explore the generation of other types of metadata, such as table, value, and relationship descriptions, and their impact on text-to-SQL systems. By enhancing metadata, we aim to further improve the accuracy and usability of text-to-SQL systems.
In conclusion, the creation and utilization of detailed metadata is essential for the development of text-to-SQL models and for enabling human users of database systems. This paper is a first step in demonstrating the role of LLMs as generators of such metadata.
@@
Another limitation is that our annotators were not experts in all the specific domains covered by the dataset, such as Thrombosis Prediction. Despite this, the annotators reported confidence in their ability to accurately label column descriptions across all domains, including these more challenging areas. Nevertheless, the lack of domain-specific expertise could introduce some bias or inaccuracies in the annotation process. Future work should consider incorporating domain experts for annotation, particularly for specialized fields, to further enhance the reliability of the dataset.
@@ A.2 Composition
Yes. Each column contains a corresponding gold column description, as well as an annotated difficulty rating (Easy, Medium, Hard or Very Hard). These levels reflect the annotators’ assessment of the amount of available information necessary to generate an accurate description for a column. Examples of columns of each difficulty can be found in B.3.
@@ A.6 Distribution
The dataset will be distributed under the Creative Commons Attribution-ShareAlike 4.0 International License ($C$ BY-SA 4.0).
@@ Appendix B Appendix
B.1 Annotation Procedure and Guidelines
B.1.1 Annotating the Quality of the Original BIRD-Bench Dataset Column Descriptions
@@
B.1.2 Annotating the Quality of the Improved Column Descriptions

### v3

@@ 5 Results
Self-Evident 3.83 3.55 3.26 2.98 2.97 3.24
Context-Aided 3.66 3.28 3.18 2.89 2.79 2.84
Ambiguity-Prone 3.46 2.46 2.90 2.76 2.69 2.04
Domain-Dependent 1.93 1.24 1.59 1.47 1.44 1.12
@@
Table 2: Mean quality scores for each model’s generated descriptions across the column difficulty levels in the dataset, as rated by the human evaluators.
5.1 Human Evaluation of Generated Column Descriptions
@@
For each model’s generated output, two human annotators independently assessed the quality of the column descriptions. As previously the set of annotators was different from the ones that created the gold dataset, but with the same type of expertise. The annotations were also based on a slightly different set of criteria from what was used when rating the original and improved column descriptions, presented in Section 3.1, as we now have the gold column descriptions as a reference point.
@@
The generated outputs were scored for their usefulness, using labels (4) “Perfect”, (3) “Superfluous Information”, (2) “Missing Information”, (1) “Incorrect”, and “No Description”. The main difference between “Perfect” and “Superfluous Information”, is that the latter contains additional redundant information in addition to the gold standard that is not incorrect but deemed excessive by the annotators. A “Missing information” column description has partly correct information, but information is missing compared to the gold standard. The full instructions and definitions of the labels are presented in Appendix B.7.4.
@@
To ensure the reliability of our evaluations, we measured the inter-annotator agreement for each model’s output. The agreement among annotators was above 0.60 for all models, indicating a substantial agreement between the annotators.
The results of the model evaluations are presented in Figure 2 and a breakdown of performance over the different difficulty levels can be found in Table 2. Exact numbers of the LLM performances can be found in Appendix B.5. GPT-4o demonstrates superior capability, producing perfect descriptions for 74.94% of the columns, substantially outperforming other models in this category. However, it’s important to note that the gold dataset was created using GPT-4o, which introduces a bias in favor of this model’s performance. As such, the GPT-4o results are included primarily for comparison purposes and should be interpreted cautiously. Command R+ follows with 63.47% perfect descriptions, though it also shows notable percentages of missing (17.29%) and incorrect (16.17%) information. Mixtral 22x8B, Qwen 2 72B, and CodeStral 22B strongly favor providing redundant information, with percentages ranging from 46.80% to 72.67%. This suggests these models may be over-generating or including unnecessary details in their descriptions. Mistral 7B presents a more balanced distribution across categories, with 29.65% perfect descriptions and 40.44% superfluous information, indicating a moderate performance level.
Table 2 highlights model performance across column difficulty levels. GPT-4o leads in all categories, scoring 3.83 for "Self-Evident" and 3.66 for "Context-Aided" columns, while other models like Command R+ and Mixtral 22x8B perform slightly lower. However, GPT-4o’s scores are as mentioned likely biased, as the dataset was partly created using GPT-4o itself.
As difficulty increases with the "Ambiguity-Prone" and "Domain-Dependent" columns, all models see a drop in performance. GPT-4o scores 3.46 and 1.93, respectively, but struggles alongside the other models in these categories, with Mistral 7B showing the weakest performance. These results underscore the need for better handling of complex and ambiguous columns across all models, and indicate the need for further documentation or domain knowledge.
GPT-4o was the only model to produce "No description" outputs (6.89%), abstaining when insufficient information was available, unlike other models. The difficulty split for these columns shows GPT-4o flagged 17 as "Self-Evident," 8 as "Context-Aided," 26 as "Ambiguity-Prone" and 3 as "Domain-Dependent." While it successfully identified many ambiguous columns, it also flagged some that were straightforward. Its flagging of easier column difficulties indicates the need for better ways of handling and dealing with ambiguous columns than our approach.
We also explored using GPT-4o as an automated judge to evaluate the quality of generated column descriptions. However, the results showed low alignment with human evaluations, with Cohen’s Kappa scores ranging from 0.12 to 0.25, indicating that LLM-as-a-judge was not a reliable substitute for human judgment. We included the results in Appendix B.2 and the prompt used in Appendix B.9 for future work to explore.
5.2 Impact of Column Descriptions on Text-to-SQL Performance
Included Descriptions Command R+ Mistral 7B CodeStral 22B Qwen2 72B GPT-4o Mixtral 22Bx8
No descriptions 0.2321 0.1473 0.2529 0.2894 0.3013 0.1750
BIRD-Bench descriptions 0.2419 0.1506 0.2647 0.3116 0.3134 0.2543
Gold descriptions 0.2516 0.1649 0.2738 0.3220 0.3328 0.2716
GPT-4o generated descriptions 0.2445 0.1538 0.2637 0.2602 0.3206 0.2583
Qwen2 generated descriptions 0.2675 0.1690 0.2966 0.3349 0.3678 0.2959
Table 3: Zero-shot execution accuracy scores for different types of metadata across various LLMs.
The full results of our text-to-SQL experiments are presented in Table 3. The text-to-SQL experiments reveal that including column descriptions generally improves the performance of all tested models, with the largest gains seen when using descriptions generated by Qwen2. Across the board, Qwen2-generated descriptions consistently outperformed both the gold descriptions and GPT-4o-generated descriptions, suggesting that the additional, seemingly superfluous information in Qwen2’s descriptions is actually beneficial for LLMs. This is a surprising result, as it contradicts the annotators’ initial assessment of what constitutes useful versus redundant information.
While all models demonstrated better accuracy with Qwen2 descriptions, the largest models, such as GPT-4o and Qwen2 72B, showed the most significant performance increases. For example, GPT-4o’s accuracy rose from 0.3013 without descriptions to 0.3678 with Qwen2 descriptions, indicating that larger models, with their greater capacity to understand semantics and context, may better leverage detailed metadata. In contrast, smaller models like Mistral 7B saw more modest gains, with accuracy improving by just 2% when descriptions were included (from 0.1473 without descriptions to 0.1690 with Qwen2 descriptions).
In addition to the strong performance of the synthetically generated descriptions, our manually crafted gold descriptions also outperformed the humanly annotated BIRD-Bench descriptions across all models. This indicates that more carefully curated descriptions, whether humanly refined or synthetically generated, provide greater clarity and utility for the models.
To further explore the impact of column descriptions, an ablation study detailed in Appendix B.1 shows that models gained over 20% in accuracy when provided with descriptions for completely uninformative columns, further underscoring the important role of detailed metadata in enhancing text-to-SQL performance.
@@ 6 Conclusions and Future Work
In this work, we demonstrated that natural language descriptions and metadata are crucial for improving the performance of text-to-SQL models. Our experiments show that including detailed column descriptions, whether manually refined or synthetically generated, consistently enhances query generation accuracy across a range of LLMs. However, generating accurate descriptions for columns with inherent ambiguity remains a challenge, as models unsurprisingly struggle to handle such columns effectively.
To address this, we propose a two-step approach for handling ambiguous columns: first, automatically identify and generate descriptions for columns where meaningful interpretation signals exist within the database. Then, for columns with inherent ambiguity, a domain expert should manually craft the necessary descriptions. This strategy ensures that the LLMs can handle straightforward columns autonomously while relying on expert input where ambiguity is present, although more research is needed to identify such ambiguous columns.
Our findings also suggest that models interpret and utilize detailed information differently than humans might expect. Specifically, descriptions that were initially categorized as containing "superfluous" information led to improved query generation performance across all models. This highlights the potential value of including more detailed metadata than humans might consider necessary.
Future work will focus on understanding which specific features of the Qwen2-generated descriptions contributed to the significant performance gains seen across all models. Additionally, we aim to explore what types of information are most beneficial for improving text-to-SQL performance, particularly in the context of ambiguous columns. Finally, future research will investigate the incorporation of other metadata types, such as value descriptions, synonyms, and numerical reasoning, to further enhance the capabilities of text-to-SQL systems.
@@
Another limitation is that our annotators were not experts in all the specific domains covered by the dataset, such as Thrombosis Prediction. Despite this, the annotators reported confidence in their ability to accurately label column descriptions across all domains, including these more challenging areas. Nevertheless, the lack of domain-specific expertise could introduce some bias or inaccuracies in the annotation process. Future work should consider incorporating domain experts for annotation, particularly for specialized fields, to further enhance the dataset’s reliability.
@@ A.2 Composition
Yes. Each column contains a corresponding gold column description, as well as an annotated difficulty rating (Easy, Medium, Hard or Very Hard). These levels reflect the annotators’ assessment of the amount of available information necessary to generate an accurate description for a column. Examples of columns of each difficulty can be found in B.8.
@@ A.6 Distribution
The dataset will be distributed under the Creative Commons Attribution-ShareAlike 4.0 International License (CC BY-SA 4.0).
@@ Appendix B Appendix
B.1 Ablation Study: Analyzing the Effects of Difficult Columns
To better understand the usefulness of column descriptions in low-information scenarios, we replaced all the column names in each database with uninformative column names. The first column of the first table in database A where denoted by A11, the second column by A12, and so on.
B.2 Ablation Study: Evaluating Models Using LLM-as-a-Judge
We performed an evaluation using LLM-as-a-judge, with GPT-4o as the judge. This evaluation aimed to investigate whether there exists an automated metric that could effectively replace human evaluations.
We compared the model rankings as judged by GPT-4o to those determined by human annotators to see if there was a consistent correlation. The results, as shown in Table 4, indicate low alignment between GPT-4o and human annotators, with Cohen’s Kappa scores ranging from 0.12 to 0.25.
Annotator/LLM Command R+ CodeStral 22B Mistral 7b Qwen 2 72B Mixtral 22B
Annotator 1 vs GPT-4o 0.22 0.16 0.24 0.21 0.16
Annotator 2 vs GPT-4o 0.25 0.18 0.23 0.22 0.17
Table 4: Cohen’s Kappa scores between annotators and GPT-4o LLM judge for various models.
B.3 Annotation Scores of the Original BIRD-Bench Descriptions and LLM Improved Descriptions
Descriptions Agreement Percentage Cohen’s Kappa
BIRD-Bench descriptions 0.79 0.68
LLM improved descriptions 0.82 0.61
Difficulty Annotation Agreement Percentage Cohen’s Kappa
Difficulty rating 0.53 0.31
B.4 Annotator Agreement Scores for Generated Column Descriptions
Annotator Agreement Cohen’s Kappa
Annotator 1 vs annotator 2 (Command R+) 0.60
Annotator 1 vs annotator 2 (CodeStral 22B) 0.71
Annotator 1 vs annotator 2 (Mistral 7b) 0.71
Annotator 1 vs annotator 2 (Qwen 2 72B) 0.62
Annotator 1 vs annotator 2 (Mixtral 22B) 0.63
Table 7: Cohen’s Kappa scores between human annotators for various models.
B.5 Exact Performance Metrics of the LLM Generated Column Descriptions
Quality GPT-4o Command R+ Mixtral 22x8B Qwen 2 72B CodeStral 22B Mistral 7B
Perfect 74.94 63.47 36.34 10.85 10.84 29.65
Superfluous 8.15 3.07 46.80 72.67 70.68 40.44
Missing 6.14 17.29 8.27 8.84 8.27 7.90
Incorrect 3.88 16.17 8.58 7.65 10.21 22.01
No description 6.89 0.00 0.00 0.00 0.00 0.00
Table 8: Detailed performance metrics of LLMs in generating SQL column descriptions.
B.6 Details of Improving the Existing BIRD-Bench Column Descriptions
Label (1) BIRD-Bench (2) LLM Updated (3) Manually Updated
Perfect 40.24 70.80 100
Somewhat correct 38.80 18.80 0
Incorrect 1.88 0.06 0
I can’t tell 0.81 7.27 0
No description 18.27 3.07 0
Total 100% 100% 100%
Table 9: Quality of the original BIRD-Bench column descriptions and the LLM-improved descriptions.
To improve column descriptions, we used GPT-4o, providing it with the BIRD-Bench column descriptions, corresponding schema in CREATE_TABLE statements, example data rows from the corresponding table, and instructions to generate descriptions for given columns. The model was given clear guidelines and in-context examples of the desired description format. The full prompt template can be found in Appendix B.9.
Two human annotators, who are authors of this paper, fluent in English and experts in SQL, independently assessed the original BIRD-Bench descriptions and the LLM-generated descriptions. They assigned each description to one of the following categories: Perfect, indicating that the description contains sufficient information for an unambiguous interpretation of the column; Somewhat Correct, where the description is somewhat correct, but there is room for improvement; Incorrect, if the description contains inaccuracies; No Description, when a description is missing; and I can’t tell, when the annotator is unable to determine the accuracy of the description based on the available data.
The full annotator guidelines and more detailed descriptions of the labels are presented in Appendix B.7.2.
Annotators’ assessments of the three datasets are shown in Table 9. There was a substantial level of agreement between annotators, with a Cohen’s kappa of $0.680.68$ on the original BIRD-Bench descriptions and $0.610.61$ on the improved generated descriptions. The GPT-4o descriptions showed a significant improvement over the original BIRD-Bench descriptions. While not perfect, LLM-generated descriptions can be a useful tool for improving existing column descriptions when there are already quality issues in the existing descriptions.
B.7 Annotation Procedure and Guidelines
B.7.1 Annotating the Quality of the Original BIRD-Bench Dataset Column Descriptions
@@
B.7.2 Annotating the Quality of the Improved Column Descriptions